OpenStudy (anonymous):

SAT question: If (a+b)^2 = 64 and a^2+b^2 = 40, what is the value of 4ab?

OpenStudy (owlcoffee):

use the second relation, solve out one of them, and then replace after doing that biniomal exponent.

OpenStudy (***[isuru]***):

hi, (a + b)^2 = a^2 + b^2 +2ab hence a^2 + b^2 = 40 (a + b)^2 = 40 + 2ab 64 = 40 +2ab 2ab = 24 4ab = 48 hope this will help ya out !!!

OpenStudy (anonymous):

thank you very much!

OpenStudy (***[isuru]***):

u r welcome